# Effective and Transparent RAG: Adaptive-Reward Reinforcement Learning for Decision Traceability

### v1

@@ 4.1 Datasets
We use three widely-used and challenging multi-hop QA datasets for evaluation: HotpotQA [51], 2WikiMultiHopQA [15], and MuSiQue [45]. Each QA consists of a natural language question, a final answer, and 10–20 retrieved paragraphs from Wikipedia (predefined in the public datasets). We randomly selected a portion from the whole dataset for training and testing as previous studies [42, 47, 19], statistics are shown in Table 2. Detailed dataset preprocessing steps are described in Appendix B.
@@
2WikiMultiHopQA Train 10,000 10 7 7839 3 2151 2.39
2WikiMultiHopQA Test 500 10 0 406 0 94 2.19
@@
4.2 Evaluation metrics
Following prior work [51, 8], we evaluate model performance using three metrics: normalized Exact Match (EM), F1 score (F1), and LLM-as-a-Judge (LJ) [54]. EM and F1 capture surface-level correctness, while LJ uses GPT-4o [29] (version 2024-11-20) to assess semantic correctness. The implementation details of all three metrics are provided in Appendix C.
4.3 Baselines
Prompt-based RAG models. We select open-source reasoning models of similar scale, including DeepSeek-R1-Distill-Qwen-7B and Qwen3-8B, as well as our backbone models Qwen2.5-7B-Instruct and Llama3.1-8B-Instruct. Those models answer the question with references directly [^2] For reasoning-capable models, the internal thinking process is implicitly conducted during generation in the direct setting.. We also include two prompting-based methods: SuRe [20], which enhances LLM performance via summarized retrieval, and self-ask [33], which decomposes complex questions into sub-questions and answers them iteratively. These serve as strong zero-shot and few-shot baselines.
SFT-based RAG models. We include two supervised fine-tuning baselines: SFT-direct, which directly generates the final answer, and SFT-reasoning, a reasoning-capable model that outputs structured responses. Training details for both models are provided in Appendix D.
RL-based RAG models. This category includes models trained with reinforcement learning. We first implement a Naive GRPO baseline using the original GRPO algorithm [38], where models are trained to produce <think></think> and <answer></answer> format responses with format and accuracy reward. Three SOTA RL-based reasoning frameworks are also involved: R1-Searcher [42], SimpleDeepSearcher [43], and ReSearch [3]. Each of these methods introduces a unique approach to reasoning—such as external search invocation, distillation-driven retrieval, or search-integrated generation. For consistency, we use their released checkpoints and replace their retrieved contexts with those provided by our datasets.
Table 3: Main experimental results across three datasets and three evaluation metrics. Models are grouped by training methodology. Our ARENA models are placed alongside their corresponding base models at the bottom for direct comparison. Bold indicates the best score; underline indicates the second-best. Metrics: EM = Exact Match (%), F1 = F1 score (%), LJ = LLM-as-a-Judge (%).
Model HotpotQA 2WikiMultiHopQA MuSiQue
EM F1 LJ EM F1 LJ EM F1 LJ
Prompt-based models
DeepSeek-R1-Distill-Qwen-7B 0.332 0.487 0.712 0.290 0.407 0.658 0.116 0.184 0.278
Qwen3-8B 0.582 0.719 0.768 0.652 0.727 0.784 0.336 0.397 0.394
SuRe-GPT-4o 0.490 0.695 0.742 0.490 0.607 0.632 0.194 0.307 0.324
SuRe-Qwen-7B 0.504 0.648 0.696 0.364 0.448 0.440 0.146 0.233 0.226
self-ask-GPT-4o 0.408 0.577 0.662 0.430 0.565 0.614 0.162 0.305 0.346
self-ask-Qwen-7B 0.016 0.112 0.246 0.004 0.128 0.086 0.006 0.052 0.087
SFT-based models
Qwen-7B-SFT-direct 0.498 0.628 0.662 0.540 0.609 0.620 0.168 0.254 0.214
Qwen-7B-SFT-reasoning 0.400 0.532 0.599 0.536 0.615 0.622 0.112 0.172 0.174
RL-based models
Naive-GRPO 0.534 0.662 0.732 0.620 0.684 0.720 0.332 0.435 0.420
R1-Searcher-7B 0.590 0.731 0.791 0.632 0.704 0.730 0.250 0.370 0.327
SimpleDeepSearcher-7B 0.500 0.633 0.686 0.632 0.707 0.748 0.258 0.347 0.344
ReSearch-7B 0.578 0.729 0.778 0.586 0.658 0.684 0.300 0.398 0.407
Ours (ARENA)
Llama3.1-8B-Instruct (Original) 0.528 0.676 0.738 0.398 0.478 0.484 0.248 0.370 0.324
ARENA-Llama-8B 0.552 0.703 0.780 0.622 0.710 0.735 0.358 0.445 0.466
Qwen2.5-7B-Instruct (Original) 0.484 0.628 0.660 0.334 0.424 0.412 0.252 0.354 0.306
ARENA-Qwen-7B 0.628 0.762 0.812 0.660 0.752 0.786 0.400 0.520 0.508
4.4 Implementation details
We conduct ARENA on Qwen2.5-7B-Instruct and Llama3.1-8B-Instruct. Our framework is implemented based on the Open-R1 [7]. During inference, we concatenate all retrieved paragraphs provided by the dataset in their original order. All models are trained using 8 NVIDIA A100-80G GPUs, with 7 GPUs allocated for policy optimization and 1 GPU dedicated to rollout inference via a vLLM [21] engine. More training details are introduced in Appendix E.
5 Experimental Results
5.1 Main results
As shown in Table 3, ARENA-Qwen-7B achieves the best performance across all datasets and metrics, consistently outperforming all baselines of similar scale. Compared to its backbone models, ARENA brings 10–30% improvements, with particularly strong gains on MuSiQue, highlighting its ability to handle longer and more complex reasoning chains. ARENA-Llama-8B also outperforms the original Llama3.1-8B-Instruct significantly, showing that our method is capable of distinct backbones. We also observe that RL-based models consistently outperform SFT-based models, confirming the advantage of learning through reward-guided exploration. In contrast, supervised fine-tuning fails to deliver meaningful gains, suggesting that it struggles to equip models with structured reasoning.
Table 4: Evaluation of ARENA-Qwen-7B under different training dataset combinations. “+WM” denotes training on 2WikiMultiHopQA + MuSiQue, “+HM” on HotpotQA + MuSiQue, and “+HWM” on all three datasets. Results are reported on all three test sets to assess cross-domain generalization.
Model HotpotQA 2WikiMultiHopQA MuSiQue
EM F1 LJ EM F1 LJ EM F1 LJ
ARENA-Qwen-7B + WM 0.568 0.711 0.764 0.642 0.732 0.734 0.394 0.504 0.486
ARENA-Qwen-7B + HM 0.598 0.741 0.796 0.578 0.671 0.680 0.404 0.519 0.478
ARENA-Qwen-7B + HWM 0.628 0.762 0.812 0.660 0.752 0.774 0.400 0.520 0.508
5.2 Generalization ability
The generalization ability of ARENA-Qwen-7B when trained on different subsets of the available datasets is shown in Table 4. Specifically, “+WM” and “+HM” denote training on two datasets while holding out one, allowing us to assess zero-shot transfer to an unseen domain. In both cases, the model performs strongly on the held-out dataset, indicating that the reasoning capabilities learned by ARENA generalize well across different data distributions. This supports our claim that the improvements from our method are not tied to a specific dataset, but reflect transferable multi-hop reasoning skills.
Figure 2: Ablation results comparing Naive GRPO, ARENA without KL stabilization, and full ARENA on three evaluation metrics across datasets. ARENA shows consistent improvements in output quality and interpretability.
5.3 Ablation study
We evaluate the effect of ARENA’s key components by comparing three variants: Naive GRPO (basic reward and flat output format), w/o KL stabilization (adaptive reward without stable KL estimation), and the full ARENA-Qwen-7B. As shown in Figure 2, ARENA achieves the best results across all datasets in LLM-as-a-Judge score, format score, and relevance score, demonstrating its more powerful ability in generate accurate and interpretable outputs.
Table 5: Comparison between ARENA-Qwen-7B and three strong closed-source models on HotpotQA. ARENA shows strong competitiveness despite smaller model size.
Model EM F1 LJ
GPT-4o 0.628 0.788 0.820
OpenAI-o1 0.656 0.816 0.890
DeepSeek-R1 0.618 0.781 0.836
ARENA-Qwen-7B 0.628 0.762 0.812
Table 6: Comparison of different accuracy reward formulations used during RL training. EM-based reward leads to the best performance on HotpotQA across multiple evaluation dimensions.
Judge Method EM F1 LJ
GPT-4o 0.612 0.752 0.804
F1 0.600 0.738 0.794
EM (ours) 0.628 0.762 0.812
5.4 Further analyses
ARENA Vs. SOTA Commercially-developed Models. We compare ARENA-Qwen-7B with three strong commercially-developed LLMs on HotpotQA in subsection 5.3, including GPT-4o &\&& OpenAI-o1 from OpenAI and DeepSeek-R1. Despite its smaller size, our ARENA shows competitive results across all metrics, with performance close to GPT-4o in particular. This highlights ARENA’s potential as a lightweight yet effective alternative for interpretable multi-hop reasoning.
Reward Analysis. We also test different reward formulations for the accuracy component during RL training, including EM, F1, and LLM-based judgment. As shown in subsection 5.3, using Exact Match as the reward provides the best overall results across EM, F1, and LJ evaluation. This confirms that a simple, binary EM-based reward provides a more stable and effective training signal for multi-hop QA tasks. More experimental results and analyses on 2WikiMultiHopQA and MuSiQue datasets are provided in Appendix G.
Case Study. Figure 3 illustrates a representative case. While Naive GRPO produces a valid-looking but unsupported answer, ARENA identifies the correct references and generates a transparent reasoning path. The benefit of KL stabilization is further supported by training curves (see Appendix F), where the stabilized version maintains smooth KL values and avoids extreme divergence spikes.
6 Conclusion
We propose ARENA, an Adaptive-Rewarded Evidence Navigation Agent that enhances the generator component of RAG systems via reinforcement learning with task-specific rewards. ARENA enables models to identify relevant evidence, reason in a structured and interpretable way, and produce accurate, verifiable answers. Extensive experiments across three multi-hop QA benchmarks demonstrate that ARENA significantly outperforms all open-source models of similar scale, with 10–30% gains in accuracy. It also generalizes well across datasets and backbone models, highlighting its robustness and transferability. By generating explicitly structured outputs, ARENA facilitates transparent reasoning, addressing a critical need for interpretability in real-world applications.
Limitation
Despite its effectiveness and transparent, ARENA assumes access to high-quality retrieved content and may underperform when critical evidence is missing or noisy. Its reward and KL stabilization are designed for RAG, which may unable to be direct adopted to other domains. Future work may explore retrieval-aware reward shaping and extend the framework to other tasks.
Appendix A Settings of the Pilot Experiments
We present a comprehensive comparison of multiple vanilla models under two inference settings: direct, where the model generates the answer based only on the question, and direct RAG, where the model receives both the question and all retrieved contexts as input. As shown in Table 7, RAG inputs consistently improve the performance of all models, highlighting the effectiveness of retrieval-augmented generation.
Moreover, even under the same RAG context, different LLMs exhibit large performance discrepancies. Notably, models like OpenAI-o1 and DeepSeek-R1 outperform open-source 7B models by 15–35% across multiple datasets and metrics. These results underscore the variability of RAG generator capabilities and motivate our work to improve them through reinforcement learning.
Table 7: Performance of vanilla models under two inference strategies: direct and direct RAG. All models are evaluated using identical retrieved contexts. Bold indicates the highest value per column. Metrics: EM = Exact Match (%), F1 = F1 score (%), LJ = LLM-as-a-Judge (%).
Model HotpotQA 2WikiMultiHopQA MuSiQue
EM F1 LJ EM F1 LJ EM F1 LJ
direct
Qwen2.5-7B-Instruct 0.188 0.247 0.262 0.222 0.253 0.234 0.040 0.090 0.061
Llama3.1-8B-Instruct 0.194 0.283 0.284 0.210 0.246 0.232 0.038 0.080 0.054
DeepSeek-R1-Distill-Qwen-7B 0.086 0.150 0.162 0.144 0.206 0.234 0.012 0.043 0.026
Qwen3-8B 0.248 0.344 0.367 0.220 0.263 0.264 0.030 0.092 0.060
GPT-4o 0.378 0.500 0.500 0.342 0.386 0.366 0.100 0.192 0.124
OpenAI-o1 0.510 0.667 0.726 0.530 0.634 0.684 0.292 0.404 0.392
DeepSeek-R1 0.422 0.564 0.590 0.454 0.536 0.546 0.202 0.330 0.294
direct RAG
Qwen2.5-7B-Instruct 0.484 0.628 0.660 0.334 0.424 0.412 0.252 0.354 0.306
Llama3.1-8B-Instruct 0.528 0.676 0.738 0.398 0.478 0.484 0.248 0.370 0.324
DeepSeek-R1-Distill-Qwen-7B 0.332 0.487 0.712 0.290 0.407 0.658 0.116 0.184 0.278
Qwen3-8B 0.582 0.719 0.768 0.652 0.727 0.784 0.336 0.397 0.394
GPT-4o 0.628 0.788 0.820 0.606 0.687 0.717 0.505 0.617 0.593
OpenAI-o1 0.656 0.816 0.890 0.702 0.805 0.837 0.622 0.744 0.762
DeepSeek-R1 0.618 0.781 0.836 0.721 0.801 0.838 0.559 0.702 0.703
Appendix B Dataset Preprocessing
We follow the preprocessing pipeline from PIKE-RAG [47] to format the train and dev splits for all datasets. Each example includes:
  • A natural language question.
  • A list of retrieved paragraphs from Wikipedia.
  • A gold-standard answer (often as a list of acceptable synonyms).
For training, we randomly sample 10,000 instances each from HotpotQA and 2WikiMultiHopQA, and 5,000 from MuSiQue. For testing, we sample 500 development examples from each dataset, following the setup in [42, 47, 19] to ensure comparability.
@@ Appendix B Dataset Preprocessing
Appendix C Details of Evaluation Metrics

### v2

@@ 4.1 Datasets
We train and evaluate our method on three widely-used multi-hop QA benchmarks: HotpotQA (Yang et al., 2018), 2WikiMultiHopQA(2Wiki) (Ho et al., 2020), and MuSiQue (Trivedi et al., 2022). Following prior work (Song et al., 2025a; Wang et al., 2025a; Jin et al., 2025), we randomly sample a total of 25,000 examples from the original training sets for training, 500 examples from each dataset’s test set for evaluation. Detailed statistics and preprocessing steps are provided in Appendix C.
4.2 Evaluation Metrics
We assess model performance using three standard multi-hop QA metrics (Yang et al., 2018; Gao et al., 2023): normalized Exact Match (EM), F1 score (F1), and LLM-as-a-Judge (LJ) (Zheng et al., 2023). EM and F1 measure surface-level string overlap with the gold answer, while LJ uses GPT-4o (OpenAI, 2024) (version 2024-11-20) to provide a semantic judgment of answer correctness. The implementation details of these three metrics are provided in Appendix D. In addition to these, we introduce two auxiliary evaluation metrics: Format score and Relevance score, which assess instruction adherence and citation accuracy. These scores follow the same evaluation criteria as used in the reward computation during training (see Section 3.3).
Table 2: Main experimental results across three datasets and three evaluation metrics. Models are grouped by training methodology. Bold indicates the best score; underline indicates the second-best. Metrics: EM = Exact Match(%), F1 = F1 score(%), LJ = LLM-as-a-Judge(%).
Model HotpotQA 2Wiki MuSiQue
EM F1 LJ EM F1 LJ EM F1 LJ
Prompt-based models
Llama3.1-8B-Instruct 52.8 67.6 73.8 39.8 47.8 48.4 24.8 37.0 32.4
Qwen2.5-7B-Instruct 48.4 62.8 66.0 33.4 42.4 41.2 25.2 35.4 30.6
DeepSeek-R1-Distill-Qwen-7B 33.2 48.7 71.2 29.0 40.7 65.8 11.6 18.4 27.8
Qwen3-8B 58.2 71.9 76.8 65.2 72.7 78.4 33.6 39.7 39.4
SuRe-GPT-4o 49.0 69.5 74.2 49.0 60.7 63.2 19.4 30.7 32.4
SuRe-Qwen-7B 50.4 64.8 69.6 36.4 44.8 44.0 14.6 23.3 22.6
self-ask-GPT-4o 40.8 57.7 66.2 43.0 56.5 61.4 16.2 30.5 34.6
SFT-based models
Qwen-7B-SFT-direct 49.8 62.8 66.2 54.0 60.9 62.0 16.8 25.4 21.4
Qwen-7B-SFT-reasoning 40.0 53.2 59.9 53.6 61.5 62.2 11.2 17.2 17.4
SimpleDeepSearcher-7B 50.0 63.3 68.6 63.2 70.7 74.8 25.8 34.7 34.4
RL-based models
Naive-GRPO 53.4 66.2 73.2 62.0 68.4 72.0 33.2 43.5 42.0
R1-Searcher-7B 59.0 73.1 79.1 63.2 70.4 73.0 25.0 37.0 32.7
Search-R1-7B 53.2 68.5 76.6 51.8 61.7 66.0 30.6 40.8 43.2
ZeroSearch-7B 51.8 64.6 70.6 43.6 51.7 53.2 27.0 36.4 34.2
ReSearch-7B 57.8 72.9 77.8 58.6 65.8 68.4 30.0 39.8 40.7
Ours (ARENA)
ARENA-Llama3.1-8B 55.2 70.3 78.0 62.2 71.0 73.5 35.8 44.5 46.6
ARENA-Qwen2.5-7B 62.8 76.2 81.2 66.0 75.2 78.6 40.0 52.0 50.8
4.3 Baselines
We compare our method against a comprehensive set of baselines across three categories: (1) Prompt-based models: we include Qwen2.5-7B-Instruct (Yang et al., 2024), Llama3.1-8B-Instruct (Grattafiori et al., 2024), DeepSeek-R1-Distill-Qwen-7B (Guo et al., 2025), and Qwen3-8B (Yang et al., 2025), which directly answer questions with references [^2] For reasoning-capable models, the internal thinking process is implicitly conducted during inference.. We also consider SuRe (Kim et al., 2024), which enhances performance via summarized retrieval, and self-ask (Press et al., 2023), which decomposes complex questions into sub-questions and answers them iteratively. (2) SFT-based models: we implement two variants: SFT-Direct, which directly generates the final answer, and SFT-Reasoning, which produces structured thoughts and answers. Implementation details are provided in Appendix E. We also include SimpleDeepSearcher (Sun et al., 2025b), which trains models using automatically generated reasoning and retrieval trajectories. (3) RL-based models: we implement a Naive-GRPO baseline using the original GRPO algorithm (Shao et al., 2024), where models are trained to output structured reasoning with format and accuracy rewards. In addition, we evaluate against four recent RL-based RAG frameworks: R1-Searcher (Song et al., 2025a), Search-R1 (Jin et al., 2025), ZeroSearch (Sun et al., 2025a), and ReSearch (Chen et al., 2025a), which all aim to teach the model to actively retrieve and reason through reinforcement learning. For consistency, we use their released checkpoints and replace their retrieved contexts with those provided by datasets.
4.4 Implementation details
We conduct ARENA on Qwen2.5-7B-Instruct and Llama3.1-8B-Instruct. Our framework is implemented based on the Open-R1 (Face, 2025). During inference, we concatenate all retrieved paragraphs provided by the dataset in their original order. All models are trained using 8 NVIDIA A100-80G GPUs, with 7 allocated for policy optimization and 1 dedicated to rollout inference via a vLLM (Kwon et al., 2023) engine. More training details are introduced in Appendix F.
Table 3: Evaluation of Format and Relevance scores (%) for ARENA-7B and Qwen2.5-7B-Instruct on three datasets. ARENA-7B shows consistent and substantial improvements across all datasets.
Model HotpotQA 2Wiki MuSiQue
Format Relevance Format Relevance Format Relevance
Qwen2.5-7B-Instruct 91.6 68.3 97.8 67.3 80.2 48.4
ARENA-7B 100.0 91.3 100.0 90.8 100.0 59.5
Figure 2: Evaluation results on 2Wiki under real-world local search and web search scenarios. ARENA-7B consistently achieves the best performance across all metrics and search modalities, demonstrating its superior generalization and reasoning capabilities in multi-hop question answering.
5 Experimental Results
5.1 Main Results
As shown in Table 2, ARENA-Qwen2.5-7B (ARENA-7B) achieves the best performance across all datasets and metrics, consistently outperforming all baselines of similar scale. ARENA-Llama3.1-8B also outperforms the original Llama3.1-8B-Instruct significantly, validating the effectiveness of our training method across distinct backbones.
Transparency. One key goal of ARENA is to improve the traceability and interpretability of RAG generator. We quantitatively evaluate the models before and after RL training by Format score and Relevance score, which measure instruction-following behavior and citation accuracy. As shown in Table 3, ARENA-7B significantly outperforms Qwen2.5-7B-Instruct in both metrics across all datasets. These results indicate that our method greatly enhances the transparency of the generated responses, enabling more faithful and controllable reasoning.
Stability. We not only provide a theoretical analysis showing the superiority of the $k_{2}$ estimator over the commonly used $k_{3}$ estimator, but also validate this claim empirically. As shown in the KL divergence curves during training (see Appendix G), the stabilized variant based on $k_{2}$ maintains smoother KL values and avoids extreme spikes, confirming its effectiveness in preventing degenerate updates during RL optimization.
Table 4: Evaluation of ARENA-7B under different training dataset combinations. “+WM” denotes training on 2Wiki + MuSiQue, “+HM” on HotpotQA + MuSiQue, and “+HWM” on all three datasets. Results are reported on all three test sets to assess cross-domain generalization.
Model HotpotQA 2Wiki MuSiQue
EM F1 LJ EM F1 LJ EM F1 LJ
ARENA-7B + WM 56.8 71.1 76.4 64.2 73.2 73.4 39.4 50.4 48.6
ARENA-7B + HM 59.8 74.1 79.6 57.8 67.1 68.0 40.4 51.9 47.8
ARENA-7B + HWM 62.8 76.2 81.2 66.0 75.2 77.4 40.0 52.0 50.8
Figure 3: Radar chart comparing six variants across five evaluation metrics on HotPotQA. ARENA shows consistent improvements in output quality and interpretability.
5.2 Generalization Ability
We demonstrate that our method generalizes well across unseen tasks, models, and retrieval settings.
Dataset generalization. As shown in Table 4, we evaluate ARENA-7B trained on different combinations of datasets to assess zero-shot transfer across domains. In both cases, the model performs strongly on the held-out dataset, indicating that the reasoning capabilities learned by ARENA generalize well across different data distributions.
Retrieval setting generalization. While our primary experiments use dataset-provided references, we also evaluate ARENA under realistic retrieval settings. We conduct experiments using local search from Wikipedia (Petroni et al., 2021), and web search via DuckDuckGo (Hands, 2012) and Bocha (Bochaai, ). Results in Figure 2 show that ARENA consistently outperforms baselines under all retrieval settings, demonstrating its robustness and effectiveness in real-world scenarios.
Model and task generalization. Our method also brings performance gains on smaller models (Qwen2.5-3B-Instruct), and shows strong transferability when applied to a summarization task using a similar RL approach. See Appendix H for details.
5.3 Ablation Study
We conduct an ablation study to evaluate the contribution of each component in our method. As shown in Figure 3, we compare six model variants on HotpotQA across five metrics: EM, F1, LJ, Format score, and Relevance score.
Naive-GRPO represents the basic GRPO setup. w/o KL stabilization removes our improved estimator for KL divergence. w/o bonus reward disables the bonus term in the adaptive reward design (see Appendix I for further analysis). F1 accuracy and LJ accuracy use F1 score and LLM-as-a-Judge score to compute the accuracy reward.
Results show that ARENA-7B consistently outperforms all other variants across all evaluation dimensions, demonstrating its advantage in both effectiveness (EM, F1, LJ) and transparency (Format, Relevance). These results confirm the value of each component in our design.
5.4 Further analyses
ARENA vs. Advanced Commercially-developed Models. Our experiments indicate that 7B-scale models exhibit a significant gap in RAG generation capability compared to advanced closed-source models. However, after reinforcement learning, our ARENA-7B achieves performance that is comparable to these commercially-developed models across multiple metrics (See Appendix J).
Accuracy Across Reasoning Hops. We further investigate how reasoning complexity affects performance by grouping questions based on their required number of hops. Our analysis shows that questions requiring fewer hops are consistently easier to answer and retrieve relevant evidence. See Appendix K for detailed results and discussion.
Case Study. Figure 4 illustrates a representative case while Naive GRPO produces a valid-looking but wrong answer, ARENA identifies the correct references and generates a transparent reasoning.
6 Conclusion
We propose ARENA, an Adaptive-Rewarded Evidence Navigation Agent that enhances the generator component of RAG systems via reinforcement learning with task-specific rewards. ARENA enables LLMs to explicitly link selected references, reasoning traces, and final answers. We further introduce a quantitative evaluation of citation relevance, making decision paths both auditable and measurable. We also study KL estimators for RAG training, develop stabilization strategies, and provide supporting empirical evidence of improved training stability and output quality. Extensive experiments on three multi-hop QA benchmarks show 10–30% accuracy gains over open-source models of similar scale, together with strong generalization across datasets and tasks. By combining transparent, evaluable reasoning with robust optimization, ARENA advances both interpretability and effectiveness for real-world RAG applications.
Appendix A Use of Large Language Models
We use Large Language Models to aid or polish writing.
Appendix B Theoretical Analysis of KL Estimators
Building on derivations from prior work Schulman (2020); Liu et al. (2025), we aim to show that the $k_{2}$ estimator yields more stable training than the $k_{3}$ estimator.
Notation.
Let $q\!\sim\!\mathcal{D}$ denote inputs, $o\!\sim\!\pi_{\theta}(\cdot\mid q)$ denote samples from the current policy, and $\pi_{\mathrm{ref}}$ be a fixed reference policy. Define
$r(q,o_{i})$ $=\;\frac{\pi_{\mathrm{ref}}(o_{i,<t}\mid q)}{\pi_{\theta}(o_{i,<t}\mid q)}\,.$
KL divergence and its gradient.
$\mathbb{D}_{\mathrm{KL}}\!\big(\pi_{\theta}(\cdot\mid q)\,\|\,\pi_{\mathrm{ref}}(\cdot\mid q)\big)$ $=\;\mathbb{E}_{o\sim\pi_{\theta}}\!\left[\log\frac{\pi_{\theta}(o_{i,<t}\mid q)}{\pi_{\mathrm{ref}}(o_{i,<t}\mid q)}\right]\,$,
$-\nabla_{\theta}\mathbb{D}_{\mathrm{KL}}$ $=\;-\,\mathbb{E}_{o\sim\pi_{\theta}}\!\left[\log\frac{\pi_{\theta}(o_{i,<t}\mid q)}{\pi_{\mathrm{ref}}(o_{i,<t}\mid q)}\;\nabla_{\theta}\log\pi_{\theta}(o_{i,<t}\mid q)\right]$
$=\;-\,\mathbb{E}\!\left[-\log r(q,o_{i})\;\nabla_{\theta}\log\pi_{\theta}(o_{i,<t}\mid q)\right]\,.$
$k_{2}$ estimator.
$L_{k_{2}}(\theta)$ $=\;\frac{1}{2}\,\mathbb{E}_{q,o}\!\left[\big(\log r(q,o_{i})\big)^{2}\right]\,$,
$\nabla_{\theta}L_{k_{2}}$ $=\;\mathbb{E}_{q,o}\!\left[\log r(q,o_{i})\;\nabla_{\theta}\log r(q,o_{i})\right]\,$,
$\nabla_{\theta}\log r(q,o_{i})$ $=\;\nabla_{\theta}\!\left(\log\pi_{\mathrm{ref}}(o_{i,<t}\mid q)-\log\pi_{\theta}(o_{i,<t}\mid q)\right)\;=\;-\,\nabla_{\theta}\log\pi_{\theta}(o_{i,<t}\mid q)\,$,
$-\nabla_{\theta}L_{k_{2}}$ $=\;\mathbb{E}_{q,o}\!\left[\log r(q,o_{i})\;\nabla_{\theta}\log\pi_{\theta}(o_{i,<t}\mid q)\right]\,.$
$k_{3}$ estimator.
$L_{k_{3}}(\theta)$ $=\;\mathbb{E}_{q,o}\!\left[r(q,o_{i})-\log r(q,o_{i})-1\right]\,$,
$\nabla_{\theta}L_{k_{3}}$ $=\;\mathbb{E}_{q,o}\!\left[\nabla_{\theta}r(q,o_{i})-\nabla_{\theta}\log r(q,o_{i})\right]\,$,
$\nabla_{\theta}r(q,o_{i})$ $=\;r(q,o_{i})\;\nabla_{\theta}\log r(q,o_{i})\;=\;-\,r(q,o_{i})\;\nabla_{\theta}\log\pi_{\theta}(o_{i,<t}\mid q)\,$,
$-\nabla_{\theta}L_{k_{3}}$ $=\;\mathbb{E}_{q,o}\!\left[\big(r(q,o_{i})-1\big)\;\nabla_{\theta}\log\pi_{\theta}(o_{i,<t}\mid q)\right]\,.$
Comparison.
(i) Gradient equivalence: $\nabla_{\theta}L_{k_{2}}=\mathbb{E}\!\left[\!-\,\log r\,\nabla_{\theta}\log\pi_{\theta}\right]$ matches $\nabla_{\theta}\mathbb{D}_{\mathrm{KL}}$ in expectation, whereas $\nabla_{\theta}L_{k_{3}}=\mathbb{E}\!\left[(r-1)\,\nabla_{\theta}\log\pi_{\theta}\right]\neq-\nabla_{\theta}\mathbb{D}_{\mathrm{KL}}$.
(ii) Bias under large deviation: when $\pi_{\theta}$ and $\pi_{\mathrm{ref}}$ diverge in later training (e.g., $r\!\gg\!1$ or $r\!\ll\!1$), the (r1)(r-1) weighting in $k_{3}$ induces a large bias relative to the KL gradient, yielding mis-scaled updates.
(iii) Symmetry of penalties: the $k_{2}$ loss $\tfrac{1}{2}(\log r)^{2}$ penalizes positive and negative deviations symmetrically around r=1r\!=\!1, aligning with the goal of RL regularization for LLMs; by contrast, $k_{3}$ is asymmetric and can over- or under-penalize depending on deviation direction.
Appendix C Datasets, Preprocessing, and Evaluation Setup
We follow the preprocessing pipeline from PIKE-RAG (Wang et al., 2025a) to format the train and dev splits for all datasets. Each example includes:
  • A natural language question.
  • A list of 10-20 retrieved paragraphs from Wikipedia.
  • A gold-standard answer (often as a list of acceptable synonyms).
Following prior work (Song et al., 2025a; Wang et al., 2025a; Jin et al., 2025), we randomly sample 10,000 training examples from HotpotQA, 10,000 from 2WikiMultiHopQA, and 5,000 from MuSiQue, totaling 25,000 training instances. For evaluation, we sample 500 development examples from each dataset. Dataset statistics are summarized in Table 5.
@@ Appendix C Datasets, Preprocessing, and Evaluation Setup
2Wiki Train 10,000 10 7 7839 3 2151 2.39
2Wiki Test 500 10 0 406 0 94 2.19
@@
Appendix D Details of Evaluation Metrics